# KG-FIT: Knowledge Graph Fine-Tuning Upon Open-World Knowledge

### v2

@@ 3.1 LLM-Guided Hierarchy Construction
ii. Bottom-Up Refinement: For each parent-child triple $(P,L,R)\in\mathbb{T}_{\mathcal{H}_{\text{split}}}$, the LLM is prompted to refine the triple using the prompt $\mathcal{P}_{\text{REFINE}}$ (Fig. 9 in Appendix E.2):
$(P^{\prime},L^{\prime},R^{\prime})=\text{LLM}(\mathcal{P}_{\text{REFINE}}(P,L,R)),$ (4)
where $(P^{\prime},L^{\prime},R^{\prime})$ represents the refined parent-child triple based on the following refinement actions:
  • No Update: $(P^{\prime},L^{\prime},R^{\prime})=(P,L,R)$.
  • Parent Merge: $P^{\prime}=P\cup L\cup R,L^{\prime}=\emptyset,R^{\prime}=\emptyset$.
  • Leaf Merge: $P^{\prime}=\{e_{1},\ldots,e_{p}\},L^{\prime}=\emptyset,R^{\prime}=\emptyset,\text{where }\{e_{1},\ldots,e_{p}\}=L\cup R$.
  • Include: $P^{\prime}=P\cup R,L^{\prime}=L,R^{\prime}=\emptyset$ or $P^{\prime}=P\cup L,L^{\prime}=\emptyset,R^{\prime}=R$.
This procedure iterates bottom-up until the root, resulting in a refined hierarchy $\mathcal{H}_{\text{LHR}}$. We place more details of the utilized prompts in Appendix E.2 and the refinement process in F.2.
@@ 3.2 Global Knowledge-Guided Local Knowledge Graph Fine-Tuning
Note: During fine-tuning, the time complexity per epoch is $O((|\mathcal{E}|+|\mathcal{T}|)\cdot n)$ where $|\mathcal{E}|$ is the number of entities, $|\mathcal{T}|$ is the number of triples, and $n$ is the embedding dimension. In contrast, most PLM-based methods [22, 28, 29, 23] have a time complexity of $O(|\mathcal{T}|\cdot L\cdot n_{\text{PLM}})$ per epoch during fine-tuning, where $L$ is the average sequence length and $n_{\text{PLM}}$ is the hidden dimension of the PLM. This is typically much higher than KG-FIT’s fine-tuning time complexity, as $|\mathcal{T}|\cdot L\gg(|\mathcal{E}|+|\mathcal{T}|)$.
4 Experiment
4.1 Experimental Setup
We describe our experimental setup as follows [^1] Our code and data are available at https://github.com/pat-jj/KG-FIT ..
Dataset #Ent. #Rel. #Train #Valid #Test
FB15k-237 14,541 237 272,115 17,535 20,466
YAGO3-10 123,182 37 1,079,040 5,000 5,000
PrimeKG 10,344 11 100,000 3,000 3,000
Experimental Strategy: For most PLM-based models, due to their high training cost, we use the results reported in their respective papers.. We reproduce PKGC [28] for all three datasets, SimKGC [27] and CSProm-KG [32] for PrimeKG. In addition, we evaluate the capabilities of LLM embeddings (TE-3-S/L: text-embedding-3-small/large [^2] https://platform.openai.com/docs/guides/embeddings/embedding-models) for zero-shot link prediction by ranking the cosine similarity between $(\mathbf{e}_{i}+\mathbf{r})$ and $\mathbf{e}_{j}$. For structure-based KGE models, we assess and present their best performance using optimal settings (shown in Table 12). For KG-FIT, we provide a detailed hyperparameter study in Table 13. We use OpenAI’s GPT-4o as the LLM for entity description generation and for LLM-guided hierarchy refinement. text-embedding-3-large is used for entity embedding initialization, which preserves semantics with flexible embedding slicing [45]. We set $\tau_{\text{min}}=0.15$ and $\tau_{\text{max}}=0.85$ for seed hierarchy construction. The values of $\tau_{\text{optim}}$ for FB15K-237, YAGO3-10, and PrimeKG are 0.52, 0.49, and 0.33, respectively. The statistics of the seed and LHR hierarchies are placed in Appendix F.3. For LCA, we designate the root as the grandparent (i.e., two levels above) source cluster node and set $m=5$. We set $\rho=0.5$, $\beta_{0}=1.2$, $\phi=0.4$, and use cosine distance for fine-tuning. Filtered setting [14, 19] is applied for link prediction evaluation. Hyperparameter studies and computational cost are detailed in Appendix I and H, respectively.
@@ 4.2 Results
Figure 5: How KG-FIT fixes overfitting (upper) and underfitting (lower) of structure-based models.
Effect of Constraints. We conduct an ablation study to evaluate the proposed constraints in Eq. 6 and 7, with results summarized in Table 3. This analysis underscores the importance of each constraint: (1) Hierarchical Distance Maintenance is crucial for both datasets. Its removal significantly degrades performance across all metrics, highlighting the necessity of preserving the hierarchical structure in the embedding space. (2) Semantic Anchoring proves more critical for the denser YAGO3-10 graph, where each cluster contains more entities, making it harder to distinguish between them based solely on cluster cohesion. The sparser FB15K-237 dataset is less impacted by the absence of this constraint. Similar to the semantics anchoring, the removal of (3) Inter-level Cluster Separation significantly affects the denser YAGO3-10 more than FB15K-237. Without this constraint, entities in YAGO3-10 may not be well-separated from other clusters, whereas FB15K-237 is less influenced. Interestingly, removing (4) Cluster Cohesion has a larger impact on the sparser FB15K-237 than on YAGO3-10. This difference suggests that sparse graphs rely more on the prior information provided by entity clusters, while denser graphs can learn this information more effectively from their abundant data.
Effect of Knowledge Sources. We explore the impact of the quality of LLM-refined hierarchies and pre-trained text embeddings on final performance, as illustrated in Figures 3 and 4. The results indicate that hierarchies constructed and text embeddings retrieved from more advanced LLMs consistently lead to improved performance. This finding underscores KG-FIT’s capacity to leverage and evolve with ongoing advancements in LLMs, effectively utilizing the increasingly comprehensive entity knowledge captured by these models.
Table 4: Model efficiency on PrimeKG. T/Ep and Inf denote training time/epoch and inference time. Our KG-FIT achieves 12 times the training speed of the best PLM-based baseline.
Method LM T/Ep Inf
KG-BERT RoBERTa 170m 2900m
PKGC RoBERTa 190m 50m
TagReal LUKE 190m 50m
StAR RoBERTa 125m 30m
KG-S2S T5 30m 110m
SimKGC BERT 20m 0.5m
CSProm-KG BERT 15m 0.2m
KG-FIT (ours) Any LLM 1.2m 0.1m
Structure-based 0.2m 0.1m
@@ 5 Conclusion
We introduced KG-FIT, a novel framework that enhances knowledge graph (KG) embeddings by integrating open-world entity knowledge from Large Language Models (LLMs). KG-FIT effectively combines the knowledge from LLM and KG to preserve both global and local semantics, achieving state-of-the-art link prediction performance on benchmark datasets. It shows significant improvements in accuracy compared to the base models. Notably, KG-FIT can seamlessly integrate knowledge from any LLM, enabling it to evolve with ongoing advancements in language models. Future work will explore incorporating LLM-generated summaries of KG triples in training set as entity descriptions, further enhancing the embedding quality. We discuss potential downstream applications of KG-FIT in Appendix J. Ethics, broader impacts, and limitations are presented in Appendix A.
Appendix A Ethics, Broader Impacts, and Limitations
A.1 Limitations
(1) Although KG-FIT outperforms state-of-the-art PLM-based models on the FB15K-237, YAGO3-10, and PrimeKG datasets, it does not outperform pure PLM-based methods on a lexical dataset WN18RR, as shown in Table 4. This limitation is discussed with details in Appendix C.
(2) KG-FIT is sensitive to the quality of the constructed hierarchy, especially the seed hierarchy. To mitigate this sensitivity, we propose automatically selecting the optimal binary tree based on the silhouette score. However, if the initial clustering is suboptimal, it may lead to a lower-quality hierarchy that impacts KG-FIT’s performance.
A.2 Ethics and Broader Impacts
@@ Appendix I Hyperparameter Study
$(P,L,R)$ Parent-child triple (sub-tree) in the hierarchy

### v3

@@ 3.1 LLM-Guided Hierarchy Construction
ii. Bottom-Up Refinement: In the bottom-up refinement phase, the LLM iteratively refines the intermediate hierarchy $\mathcal{H}_{\text{split}}$ produced by the cluster splitting step. The refinement process starts from the leaf level of the hierarchy and progresses upwards, considering each parent-child triple $(P_{*},P_{l},P_{r})$, where $P_{*}$ represents the grandparent cluster, and $P_{l}$ and $P_{r}$ represent the left and right child clusters of $P_{*}$, respectively. Let $\{C^{l}_{1},C^{l}_{2},C^{l}_{3},\ldots,C^{l}_{|P_{l}|}\}$ denote the children of $P_{l}$, and $\{C^{r}_{1},C^{r}_{2},C^{r}_{3},\ldots,C^{r}_{|P_{r}|}\}$ denote the children of $P_{r}$.
For each parent-child triple, the LLM is prompted with $\mathcal{P}_{\text{REFINE}}(P_{*},P_{l},P_{r})$ (Fig. 9 in Appendix E.2), which provides the names and entities of the grandparent and child clusters. The LLM then suggests a refinement action to update the triple based on its understanding of the relationships between the clusters. The refinement options are:
  1. 1.
    NO UPDATE: The triple remains unchanged, i.e., $(P_{*}^{\prime},P_{l}^{\prime},P_{r}^{\prime})=(P_{*},P_{l},P_{r})$.
  2. 2.
    PARENT MERGE: All the children of $P_{l}$ and $P_{r}$ are merged into the grandparent cluster $P_{*}$, resulting in $P_{*}^{\prime}=\{C^{l}_{1},C^{l}_{2},C^{l}_{3},\ldots,C^{l}_{|P_{l}|},C^{r}_{1},C^{r}_{2},C^{r}_{3},\ldots,C^{r}_{|P_{r}|}\}$. The original child clusters $P_{l}$ and $P_{r}$ are removed from the hierarchy.
  3. 3.
    LEAF MERGE: $P_{*}^{\prime}=\{e_{1},e_{2},\ldots,e_{p}\},P_{l}^{\prime}=\emptyset,P_{r}^{\prime}=\emptyset,\text{where }\{e_{1},e_{2},\ldots,e_{p}\}=P_{l}\cup P_{r}$.
  4. 4.
    INCLUDE: One of the child clusters is absorbed into the other, while the grandparent cluster remains unchanged. This can happen in two ways:
    • $P_{*}^{\prime}=P_{l}^{\prime}=\{C^{l}_{1},C^{l}_{2},C^{l}_{3},\ldots,C^{l}_{|P_{l}|},P_{r}\}$, and $P_{r}^{\prime}=\emptyset$, or
    • $P_{*}^{\prime}=P_{r}^{\prime}=\{P_{l},C^{r}_{1},C^{r}_{2},C^{r}_{3},\ldots,C^{r}_{|P_{r}|}\}$, and $P_{l}^{\prime}=\emptyset$.
The LLM determines the most appropriate refinement action based on the semantic similarity and hierarchical relationships between the clusters. The refinement process continues bottom-up, iteratively updating the triples until the root of the hierarchy is reached. The resulting refined hierarchy is denoted as $\mathcal{H}_{\text{LHR}}$. We place more details of the process in Appendix E.2 and F.2.
@@ 3.2 Global Knowledge-Guided Local Knowledge Graph Fine-Tuning
Note: During fine-tuning, the time complexity per epoch is $O((|\mathcal{E}|+|\mathcal{T}|)\cdot n)$ where $|\mathcal{E}|$ is the number of entities, $|\mathcal{T}|$ is the number of triples, and $n$ is the embedding dimension. In contrast, classic PLM-based methods [22, 28, 29, 23] have a time complexity of $O(|\mathcal{T}|\cdot L\cdot n_{\text{PLM}})$ per epoch during fine-tuning, where $L$ is the average sequence length and $n_{\text{PLM}}$ is the hidden dimension of the PLM. This is typically much higher than KG-FIT’s fine-tuning time complexity, as $|\mathcal{T}|\cdot L\gg(|\mathcal{E}|+|\mathcal{T}|)$.
@@
4 Experiment
4.1 Experimental Setup
We describe our experimental setup as follows.
Dataset #Ent. #Rel. #Train #Valid #Test
FB15k-237 14,541 237 272,115 17,535 20,466
YAGO3-10 123,182 37 1,079,040 5,000 5,000
PrimeKG 10,344 11 100,000 3,000 3,000
Experimental Strategy: For most PLM-based models, due to their high training cost, we use the results reported in their respective papers.. We reproduce PKGC [28] for all three datasets, SimKGC [27] and CSProm-KG [32] for PrimeKG. In addition, we evaluate the capabilities of LLM embeddings (TE-3-S/L: text-embedding-3-small/large) for zero-shot link prediction by ranking the cosine similarity between $(\mathbf{e}_{i}+\mathbf{r})$ and $\mathbf{e}_{j}$. For structure-based KGE models, we assess and present their best performance using optimal settings (shown in Table 12). For KG-FIT, we provide a detailed hyperparameter study in Table 13. We use OpenAI’s GPT-4o as the LLM for entity description generation and for LLM-guided hierarchy refinement. text-embedding-3-large is used for entity embedding initialization, which preserves semantics with flexible embedding slicing [45]. We set $\tau_{\text{min}}=0.15$ and $\tau_{\text{max}}=0.85$ for seed hierarchy construction. The values of $\tau_{\text{optim}}$ for FB15K-237, YAGO3-10, and PrimeKG are 0.52, 0.49, and 0.33, respectively. The statistics of the seed and LHR hierarchies are placed in Appendix F.3. For LCA, we designate the root as the grandparent (i.e., two levels above) source cluster node and set $m=5$. We set $\rho=0.5$, $\beta_{0}=1.2$, $\phi=0.4$, and use cosine distance for fine-tuning. Filtered setting [14, 19] is applied for link prediction evaluation. Hyperparameter studies and computational cost are detailed in Appendix I and H, respectively.
@@ 4.2 Results
Figure 3: KG-FIT can mitigate overfitting (upper) and underfitting (lower) of structure-based models.
[…]
Figure 3 further illustrates the robustness of KG-FIT, showing superior validation performance across training steps compared to the corresponding base models, indicating its ability to fix both overfitting and underfitting issues of some structure-based KG embedding models.
Effect of Constraints. We conduct an ablation study to evaluate the proposed constraints in Eq. 5 and 6, with results summarized in Table 4. This analysis underscores the importance of each constraint: (1) Hierarchical Distance Maintenance is crucial for both datasets. Its removal significantly degrades performance across all metrics, highlighting the necessity of preserving the hierarchical structure in the embedding space. (2) Semantic Anchoring proves more critical for the denser YAGO3-10 graph, where each cluster contains more entities, making it harder to distinguish between them based solely on cluster cohesion. The sparser FB15K-237 dataset is less impacted by the absence of this constraint. Similar to the semantics anchoring, the removal of (3) Inter-level Cluster Separation significantly affects the denser YAGO3-10 more than FB15K-237. Without this constraint, entities in YAGO3-10 may not be well-separated from other clusters, whereas FB15K-237 is less influenced. Interestingly, removing (4) Cluster Cohesion has a larger impact on the sparser FB15K-237 than on YAGO3-10. This difference suggests that sparse graphs rely more on the prior information provided by entity clusters, while denser graphs can learn this information more effectively from their abundant data.
Table 3: Model efficiency on PrimeKG. T/Ep and Inf denote training time/epoch and inference time. KG-FIT outperforms all the PLM-based models.
Method LM T/Ep Inf
KG-BERT RoBERTa 170m 2900m
PKGC RoBERTa 190m 50m
TagReal LUKE 190m 50m
StAR RoBERTa 125m 30m
KG-S2S T5 30m 110m
SimKGC BERT 20m 0.5m
CSProm-KG BERT 15m 0.2m
KG-FIT (ours) Any LLM 1.2m 0.1m
Structure-based 0.2m 0.1m
Effect of Knowledge Sources. We explore the impact of the quality of LLM-refined hierarchies and pre-trained text embeddings on final performance, as illustrated in Figures 4 and 5. The results indicate that hierarchies constructed and text embeddings retrieved from more advanced LLMs consistently lead to improved performance. This finding underscores KG-FIT’s capacity to leverage and evolve with ongoing advancements in LLMs, effectively utilizing the increasingly comprehensive entity knowledge captured by these models.
@@ 5 Conclusion
We introduced KG-FIT, a novel framework that enhances knowledge graph (KG) embeddings by integrating open-world entity knowledge from Large Language Models (LLMs). KG-FIT effectively combines the knowledge from LLM and KG to preserve both global and local semantics, achieving state-of-the-art link prediction performance on benchmark datasets. It shows significant improvements in accuracy compared to the base models. Notably, KG-FIT can seamlessly integrate knowledge from any LLM, enabling it to evolve with ongoing advancements in language models. Future work will explore incorporating LLM-generated summaries of KG triples in training set as entity descriptions, further enhancing the embedding quality. We discuss potential downstream applications of KG-FIT in Appendix J. Ethics and broader impacts are presented in Appendix A. Our code and data are available at https://github.com/pat-jj/KG-FIT.
6 Limitations
(1) Although KG-FIT outperforms state-of-the-art PLM-based models on the FB15K-237, YAGO3-10, and PrimeKG datasets, it does not outperform pure PLM-based methods on a lexical dataset WN18RR, as shown in Table 5. This limitation is discussed with details in Appendix C. As a future work, we will explore the integration of contrastive learning into the KG-FIT framework to enhance its capability to capture semantic relationships more effectively.
(2) KG-FIT’s performance is influenced by the quality of the constructed hierarchy, particularly the seed hierarchy. To address this, we propose an automatic selection of the optimal binary tree based on the silhouette score. However, if the initial clustering is suboptimal, it may result in a lower-quality hierarchy that affects KG-FIT’s performance. Additionally, the bottom-up refinement process in our proposed LHR approach updates each parent-child triple with only a single operation, which prioritizes efficiency and simplicity over performance. In future work, we plan to explore cost-efficient methods for refining the hierarchy that integrate multiple operations for each triple update, striking a better balance between efficiency and performance.
7 Acknowledgement
The research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Ethics and Broader Impacts
@@ Appendix I Hyperparameter Study
$(P_{*},P_{l},P_{r})$ Parent-child triple (sub-tree) in the hierarchy
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Section 4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: the paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Please refer to Section 4 and Appendix I to reproduce our results. We also release all the code and data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release our code and data in the uploaded zipped supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 4 and Appendix I
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Section 4.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix H.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics. The research conducted in this paper conform with it.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: In Appendix A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Section 4 and Appendix B.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: In this paper, we provide a subset of PrimeKG, which is described in Appendix B, and we do not release new assets. We shared an anonymized URL to our code and data.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.